OpenStudy (anonymous):

Johan found that the equation –2|8 – x| – 6 = –12 had two possible solutions: x = 5 and x = –11. Which explains whether his solutions are correct? He is correct because both solutions satisfy the equation. He is not correct because he made a sign error. He is not correct because there are no solutions. He is not correct because there is only one solution: x = 5.

OpenStudy (dtan5457):

He is not correct.

OpenStudy (dtan5457):

\[\frac{ -2\left| 8-x \right|=6 }{ -2 }\] Divide by -2. \[\left| 8-x \right|=3\]

OpenStudy (dtan5457):

Now since it is absolute value, this can equal 8-x=3 8-x=-3 x=5,11

OpenStudy (dtan5457):

He put 5,-11, which is wrong